The perimeter of a rectangle is 62inches, the length exceeds the width by 23inches. Find teh length and width?
Answer by avani(25)

The perimeter of a rectangle is 62inches, the length exceeds the width by 23inches. Find teh length and width?
perimeter of a rectangle= 2(L) + 2(W)
Let length (L) = x+23 inches
Width (W) = 23 inches
Perimeter of a rectangle(P) = 62inches
sub. in the formula
perimeter of a rectangle= 2(L) + 2(W)
62 = 2(x+23) + 2(23)
62 = 2x+46+46
62=2x+92
62-92=2x
-30/2=x
-15=x
As a result, length (L) = x+23 inches = -15+23 = 8 inches and Width (W) = 23 inches
Check:
perimeter of a rectangle= 2(L) + 2(W)=2(8)+2(23)=62 inches